OpenStudy (anonymous):

Need help with a quadratic equation. x^2+2x-3=0.

OpenStudy (anonymous):

x=1,-3

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

I can show you. Have you learned how to factor quadratics or know the quadratic formula?

OpenStudy (anonymous):

I'm just learning how to factor quadratics.

OpenStudy (anonymous):

watch the two lectures I posted. you will understand all you need to

OpenStudy (anonymous):

factor, then solve the factors. \[x^2 + 2x - 3 = 0\] factors to (x+3)(x-1)=0 Then you just solve independently x +3 = 0 and x-1 = 0 because if either terms (x+3) or (x-1) = 0 the whole thing equals zero. Your two answers are then x = -3 and x = 1.

OpenStudy (anonymous):

the book I have shows an example that has me getting a square root and solving that way which has me coming up with the square root of 7.

OpenStudy (anonymous):

You can ALWAYS use the quadratic formula...

OpenStudy (anonymous):

@bryanj, you're method makes more sense to me. Can all "solve by completing the square" problems be solved using the quadratic formula?

OpenStudy (anonymous):

you did some calculation error in employing the quadratic formula which involves a square root

OpenStudy (anonymous):

yes, ALL quadratics can be solved by quadratic formula. but completing the squares is more elegant in some cases.

OpenStudy (anonymous):

Ok, so I have a problem that has the answer -2-3(sq. root symbol on the 3) and -2+3(sq. root symbol on the 3) the problem is v^2+4v+1=0 Can the quadratic formula be applied the same way or is it set up differently?

OpenStudy (anonymous):

quadratic formula can be used in this case, where a = 1, b = 4 and c =1 if you take those values for a b and c and plug them into the quadratic formula you should end up with those answers

OpenStudy (anonymous):

yes, the quadratic can be applied in the same way. you already have the roots.

OpenStudy (anonymous):

ok I think I'm getting it now...I'll try a few on my own and if I need more help, I'll beb ack to this great site!

OpenStudy (anonymous):

good. do that.